40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Mar 24 05:05:09 2015, ericq, Update, LSC, AO Path engaged CARM_DARM_AOengaged_23March2015.pdfloops.png
    Reply  Tue Mar 24 18:22:11 2015, ericq, Update, LSC, AO Path engaged IMC_ao_Mar242015.pngIMC_olgs_Mar242015.png
    Reply  Tue Mar 24 18:47:10 2015, ericq, Update, LSC, AO Path engaged 
Message ID: 11168     Entry time: Tue Mar 24 18:47:10 2015     In reply to: 11163
Author: ericq 
Type: Update 
Category: LSC 
Subject: AO Path engaged 

Jenne has more detailed notes about how things went down last night, but I figure I should write about how we got the AO path stably up. 

As the carm_cm_up script stood after Jenne and Den's work last week, the CARM loop looked like the gold trace in the loop shape plot I posted in the previous elog. The phase bubble was clearly enlarged by the AO path, but there was some bad crossover instability brewing at 400 Hz. This was evident as a large noise peak, and would lead to lock loss if we tried to increase the overall CARM gain.

Quote:

 

As with our single arm CM board locking adventures, it was useful to have a filter that made the digital loop shape steeper around the crossover region, so that the 1/f AO+cavity pole shape played nice with the digital slope. As in the single arm trials, this effectively meant undoing the cavity pole compensating zero with a corresponding pole, letting the physical cavity pole do the steepening. This is only possible once the AO path has bestowed some phase upon you. A zero at a somewhat higher frequency (500Hz) gives the digital loop back some phase, which is neccesary to stay locked when the loop has only a few hundred Hz UGF, and the digital phase still matters. This gives us the purple trace. 

This provided us with a loop shape that could smoothly be ramped up in overall gain towards UGFs of multiple kHz (red trace). At this point we could reliably turn on the first boost, which will help in transitioning the PRMI to 1f signals (green trace). We didn't want to ramp it up too much, as we saw that the phase bubble likely ended not much higher than 100kHz, and the OLG magnitude was flattening pretty clearly around 40kHz. While we could turn on a super boost, it didn't look too nice, as we would have to stay at low phase margin to avoid bad gain peaking (blue trace).

As could be seen in the noise spectra that Jenne showed, you can see the violin notches in the CARM noise. This means we are injecting the digital loop noise all over the place. We attempted rolling off the digital loop (by undoing the zero at 500Hz), but found this made the gain at ~200Hz crash down, almost becoming unstable. We likely haven't positioned the crossover frequency in the ideal place for doing this. 

We didn't really give the interferometer any time to see how the long term stability was, since we wanted to poke around and measure as much as we could. While not every attempt would get us all the way there, the current carm_cm_up's success rate at achieving multi-kHz CARM bandwidth was pretty good (probably more than 50%) and the whole thing is still pretty snappy. 

ELOG V3.1.3-